OpenStudy (anonymous):

Two different plants grow each year at different rates, which are represented by the functions f(x) = 4x and g(x) = 5x + 2. What is the first year the f(x) height is greater than the g(x) height? A. Year 0 B.Year 1 C.Year 2 D.Year 3 Can i get some help plz

OpenStudy (anonymous):

@ali2x2

OpenStudy (anonymous):

@Michele_Laino can u help me plz

OpenStudy (michele_laino):

you have to substitute x=0, 1,2,3 then you have to compare the values f(x) and g(x) for each year. For example, if x=0, then we have: f(0) = 4*0 =0, and g(0)= 5*0+2=2, so f(0) < g(0) now, please do the same for other values of x

OpenStudy (anonymous):

so pu 4 in for f in the problem and 5 in for g

OpenStudy (anonymous):

u there

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

question: are f(x) and g(x) heights or rates?

OpenStudy (anonymous):

rates

OpenStudy (michele_laino):

so, are you familiar with differential calculus?

OpenStudy (anonymous):

no i am not ?

OpenStudy (michele_laino):

if f(x) is a rate, then the corresponding height h, is given solving this differential equation: \[\Large \frac{{dh}}{{dx}} = 4x\]

OpenStudy (anonymous):

ok i am trying to understand what is going on but for my first answer i put B. am i right

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

my reasoning is: height h1 for f(x) is: \[\Large {h_1}\left( x \right) = 2{x^2}\] whereas height h2 for g(x) is: \[\Large {h_2}\left( x \right) = \frac{{5{x^2}}}{2} + 2x\] and we have h2>h1, if x>0, whereas h2=h1 if x=0

OpenStudy (michele_laino):

please wait I ask to another helper

OpenStudy (anonymous):

so A

OpenStudy (michele_laino):

@Astrophysics please help

OpenStudy (michele_laino):

@vera_ewing @nincompoop please help

OpenStudy (anonymous):

so what i need to do

OpenStudy (astrophysics):

This doesn't really make sense, are the functions \[f(x)=4x~~and~~g(x) = 5x+2~~~~or~~~~f(x)=4^x~~~~and~~~g(x) = 5x+2\]

OpenStudy (anonymous):

Two different plants grow each year at different rates, which are represented by the functions f(x) = 4x and g(x) = 5x + 2. What is the first year the f(x) height is greater than the g(x) height? this is the problem

OpenStudy (astrophysics):

What are the functions...from what i showed you

OpenStudy (astrophysics):

is it to the power of x or just times x...

OpenStudy (anonymous):

just times

OpenStudy (astrophysics):

So it's \[f(x) = 4x\] and \[g(x) = 5x+2\]

OpenStudy (anonymous):

i am so confused tho

OpenStudy (astrophysics):

Can you please take an image of your question and post it

OpenStudy (anonymous):

i dont know how to do thatfrom computer

OpenStudy (anonymous):

Two different plants grow each year at different rates, which are represented by the functions f(x) = 4x and g(x) = 5x + 2. What is the first year the f(x) height is greater than the g(x) height? Year 0 Year 1 Year 2 Year 3

OpenStudy (astrophysics):

Ok all I am asking you is, what are the functions they gave you in the problem f(x), if it's to the power of x or is just 4x...

OpenStudy (astrophysics):

and same for g(x)

OpenStudy (astrophysics):